ABC Company makes 40,000 units per year of a part it uses in the
products it manufactures. The per unit product cost of this part
is shown below:

direct materials ..............  $15.30

direct labor ..................   27.40

variable overhead .............    2.10

fixed overhead ................   24.70

total .........................  $69.50

An outside supplier has offered to sell ABC Company 40,000 units
of this part a year for $66.10 per unit. If ABC Company accepts
this offer, the facilities now being used to make this part could
be used to make more units of a product that is in high demand.
The additional contribution margin that could be earned on this
other product would be $100,000 per year.

If ABC Company accepts the outside supplier's offer, $21.90 of the
fixed overhead cost being applied to the part would be eliminated.
The remaining amount would continue to be incurred and would be
allocated to the company's remaining products.

Calculate the selling price per unit charged by the outside
supplier that would make ABC Company economically indifferent
between making and buying the part
. Enter your answer with two
places after the decimal point (i.e., $78.90).
Answer #1
The cost which is relevant for the decision making of either to buy or purchase is below:
Direct material=15.3
Direct labor=27.4
Variable overhead=2.1
Avoidable overhead=24.7-21.8=2.9 (Since 21.8 is anyways will be incurred)
avoidable cost=47.7
Total avoidable cost =47.7*40000=1908000
Profit from use of freed capacity=100000
Selling price per unit charged by supplier=(1908000+100000)/40000=50.2